### v1

# Modeling Language Evolution Using a Spin Glass Approach

Abstract
The evolution of natural languages poses a riddle to any theoretical perspective based on efficiency considerations. If languages are already optimally effective means of organization and communication of thought, why do they change? And if they are driven to become optimally effective in the future, why do they change so slowly, and why do they diversify, rather than converge towards an optimum? We look here at the hypothesis that disorder, rather than efficiency, may play a dominant role.
Most traditional approaches to study diachronic language dynamics emphasize lexical data, but it would seem that a crucial contribution to the effectiveness of a thought-coding device is given by its core structure, its syntax. Based on the reduction of syntax to a set of binary syntactic parameters, we introduce here a model of natural language change in which diachronic dynamics are mediated by disordered interactions between/among parameters, even in the idealized limit of identical external inputs. We show in which region of ‘phase space’ such dynamics show the glassy features that are observed in natural language across time. In particular, binary syntactic vectors remain trapped in glassy metastable (tendentially stable) states when the degree of asymmetry in the disordered interactions is below a critical value, consistent with studies of spin glasses with asymmetric interactions. We further show that an added Hopfield-type memory term, would indeed, if strong enough, stabilize syntactic configurations even above the critical value, but losing the multiplicity of stable states. Finally, using a notion of linguistic distance in syntactic state space we show that a phylogenetic signal may remain among related languages, despite their gradually divergent syntax, exactly as recently pointed out for real-world languages. These statistical results appear to generalize beyond the dataset of 94 syntactic parameters across 58 languages, used in this study.
I Introduction
The study of languages is crucial for a profound comprehension of human cognition, culture, and history. Yet, the most basic observation, that natural languages are diverse, remains theoretically unexplained. As the myth of the Tower of Babel indicates, it would make sense for an effective means of communication to be common to all those who might need to share their mental states. The counter-argument that some populations would benefit from their language being unintelligible to others seems overall weak. Further, it would seem plausible that structurally different language forms would significantly differ in their efficiency, leading human groups to discard the less efficient ones. Yet, structurally very different natural languages survive and thrive to this day. Rather than converging under selective pressures towards a common standard, they even appear to diverge, albeit very slowly, as if drifting away from each other. Drift dynamics might be expected to characterize superficial language features, such as the lexicon, but less so the internal syntactic structures – which must indeed be more resistant to change [1] but still in the direction of increased diversity.
For two centuries, the study of language evolution has in fact largely relied on the form of lexical items [2]. Recently, Bayesian phylogenetic methods have been applied to core vocabulary [3], and geographic data has been integrated to model, e.g., the spread of Indo-European languages [4]. Historical constraints have been used to evaluate hypotheses on their place of origin [5], and automatic word comparison methods have enhanced reproducibility and bridged traditional approaches with machine learning [6].
Studying the lexicon, however, can contribute little to address the basic questions above, because words can change in a myriad ways, and it is not feasible to quantify and contrast any potential change in language efficiency with the number of possibilities effectively sampled, what in a physical system would be the energy-entropy balance. Syntax, the structuring of words into sentences, is much more constrained: parametric approaches claim to reduce its diversity to a simple string of binary parameter values associated with each natural language [7]. The number of such parameters, initially thought to be few [8], has recently grown into the hundreds [9], but they still span a much more limited set than the one spanned by lexical items [10].
In recent decades, in fact, phylogenetic linguistics has advanced through the application of quantitative methods to generative syntax and the comparison of parameter values [11, 12, 13, 14, 15, 16, 17, 9]. It has been shown that, contrary to long-held assumptions, syntactic features —- when analyzed within the generative biolinguistic framework —- can offer valuable insights into the long-term evolution of languages [17]. This approach reveals that syntax can carry a historical signal comparable to that of traditional etymological methods, but with sufficient verticality and greater chronological depth [17]. Recent research [9] has proposed that the integration of the Parametric Comparison Method (PCM) [14, 15, 17] with theories of possible languages [10] can elucidate phylogenetic relationships between language families. Following this line one may thus be able to extend the temporal scope of traditional comparative methods, offering a promising solution to long standing challenges in taxonomic linguistics [9].
Chomsky’s original Principles &\&& Parameters model describes grammatical diversity by attributing cross-linguistic variation to binary syntactic parameters. It posits a universal and innate syntactic structure, with differences among languages arising from distinct parameter settings [18, 7], which are rapidly acquired during language learning by each individual. Language acquisition, in this view, thus involves selecting from a finite set of options [19]. More recent approaches within the minimalist program question the innateness of parameters, suggesting instead that they emerge dynamically through cognitive processes [20]. Still, aside from the innateness debate, linguistic variation is seen as governed by structured binary choices, offering a systematic account of phenomena such as variability in subject or object omission or word order.
Syntactic parameter values cannot all be set, each independently of the others; parameters are interconnected through a hierarchical network of asymmetric implicational dependencies, exhibiting a near-logical structure shaped by functional constraints. For example, parameter B may be defined only if parameter A takes value +1. Seen the other way around, through the implications the value of one parameter may condition others. These can be regarded as strong, unidirectional interactions among parameters, that proceed always downstream from the top parameters, those that can be set independently of any other, toward the bottom of the hierarchy, occupied by parameters that can vary only subject to many conditions. In a network model, such dependencies may be formalized through polysynaptic weights, as detailed below.
Unlike biological evolution, where adaptive pressures can usually be identified, that drive living systems toward greater efficiency (or fitness), syntactic change does not appear to follow such a straightforward route. While historical syntactic states can often be reconstructed, the evolution of syntax - lacking clear functional directions - remains both constrained and unpredictable. Its trajectories are marked by two seemingly contradictory features: persistence, where certain structures remain stable over time, and diversification, where languages diverge into distinct syntactic systems.
The number and the complex interactions among syntactic parameters make statistical methods a reasonable option for their study. Although parameter values likely vary in subtle ways also across individuals nominally speaking the same natural language, the sets of parameter values of natural languages can be regarded as so-called metastable states. That is, they are attractive – attractive to all the individuals sharing a particular language – and not absolutely but largely stable, over centuries. The diversity expressed by the observed sets can be interpreted as stemming from the lack of any overarching order or symmetry principle, which might have imposed common values. This situation parallels glassy dynamics, in which numerous weak, disorderly interactions among components produce persistent, quasi-stable states that do not appear to conform to any ordering or symmetry principle. Therefore, disorder may play a pivotal role in syntactic evolution by enabling both diversification and persistence within language systems.
It should be emphasized that these properties appear rather trivial if considered individually: diversification – without attractor dynamics – can be explained away by different external inputs, while persistence – if diachronic changes were negligible – is obviously a generic property of static systems. To analyze, however, how these properties can coexist and interact, it seems necessary to go beyond linear reasoning and adopt a quantitative model.
In the following, in Section II, we introduce a novel model of syntactic change inspired by glassy systems. This model combines strong, explicitly modeled asymmetric interactions that tendentially produce implicational relations, with a multitude of weak, statistically modeled pairwise interactions, expressing disorder in the exact way parameters influence each other. The model can be used to investigate the overall impact of these effectively random couplings on system dynamics. Section III presents a sharp transition obtained by slightly varying an angle defining the average asymmetry in the disordered interactions, and locates the transition point. In Section IV, we incorporate a Hebbian term as an additional symmetric interaction weight between parameters, and examine its effect on evolutionary dynamics. Finally, in Section V, we focus on language distances, analyzing the ‘historical signal’, which reflects linguistic relatedness.
Figure 1: Symmetric random couplings can block syntax evolution. Simulations illustrating parameter fluidity as a function of implication strength $\zeta$ and of the degree of symmetry $\phi$ in the random interactions. The black datapoints are the number of flipping parameters as a function of time for $\zeta\to\infty$, i.e., with implications alone. Syntactic structures then continue to evolve, with 20–30 parameters flipping at each time step, indicating fluid dynamics. When adding random interactions, the strength $\zeta=0.01,0.1,1,10,100,1000$ of the implications is denoted by progressively darker colors. With random symmetric connections, ($\phi=0^{\circ}$, panel a), after the first few time steps, parameters stop changing, whatever the value of $\zeta$. Eventually, after more time steps, the same freezing occurs in panel (b), with $\phi=30^{\circ}$; with more asymmetry, instead, i.e., $\phi=60^{\circ}$, (c) and $\phi=90^{\circ}$, (d), fluidity is maintained throughout, with the parameters continuing to flip over time, actually the more the weaker are the implications. Each of 58 natural languages is used as a starting point for 10 simulations, for the same set of weak interaction weights; and then averages are taken over 100 distinct sets of weights. Therefore, the curves shown are averages over a total of 58,000 data points.
II Modeling Syntactic Evolution
This study is based on simulating a simple model of the evolution of syntax over time. The model assumes that parameters can change value (or flip) under the influence of other parameters, and with negligible external inputs. Language change therefore arises from the attempt, by the parameters, to coordinate their values, in what is effectively a recurrent network. The crucial assumption is that complete coordination may never be possible, given the lack of intentional design: a large array of disorderly constraints tends to leave those mutually incompatible, to some extent, and prone to change. The assumption of negligible external inputs does not imply that we believe them to be irrelevant to real syntax dynamics, of course; it is simply a methodological choice adopted with models of other complex systems addressing the same critical issue [21]: do they allow for a multiplicity of attractor states even when isolated from external inputs?
Internal influences are of two distinct types: implications, which explicitly reproduce linguistically derived conditions on parameter values, and are inherently asymmetric; and random couplings, which represent all presumable interactions among parameters that have not yet been analyzed explicitly, and which therefore we treat statistically. They are randomly assigned in this statistical sense, but we assume them to be the same across all languages, and keep them fixed through their evolution. Implications may involve multiple conditions, e.g. parameter $i$ can be defined and hence may take its non-default value only if either parameter $j$ or parameter $k$ is set to +1. Their overall strength (relative to the random couplings) is scaled by a coefficient $\zeta$, which implies that for $\zeta\to\infty$ implications become absolute rules, while for $\zeta\to 0$ they reduce to weak biases. The random couplings, instead, reflect the assumption that all parameters, implemented within one brain, must perforce interact in a multitude of ways not individually analyzable: to hope that they could be isolated from each other appears to be wishful thinking. In this study, we consider only the simplest interactions, i.e. pairwise ones, postponing more complex variants to future analyses. We introduce an angle $\phi$ that statistically quantifies the degree of symmetry of such random couplings, from $\phi=0^{\circ}$ (fully symmetric couplings, e.g., from parameter $j$ to parameter $i$ and viceversa the two coupling coefficients would be equal, $W_{ij}=W_{ji}$); to $\phi=90^{\circ}$ (anti-symmetric couplings, i.e., they would be opposite, $W_{ij}=-W_{ji}$).
The model is applied to the most extensive available corpus of parametric syntax: the PCM dataset put together by Longobardi and collaborators [14, 9]. It includes $L=58$ languages from various families, each characterized by $N=94$ syntactic parameters relative to the domain of nominal syntax. Each of the $58\times 94$ parameter values, denoted as $\{S_{i}^{\mu}\}$, with $i=1,\ldots,N$ and $\mu=1,\ldots,L$, is either set to $\{S_{i}^{\mu}\}=-1$ (default) or to $\{S_{i}^{\mu}\}=+1$ (the marked value). Note that our notation aims to bridge with the physics literature on spin systems, hence binary values are identified with Spins [22]; and with that on associative memories, hence distinct language configurations are indexed by the symbol $\mu$ [21]. We refrain from using the symbol $p$ for either the number of parameters considered or, as in the associative memory literature, the number of distinct patterns or configurations of values they are supposed to be able to express, and use $N$ and $L$ instead, respectively.
In the PCM database, many parameters can also take the value 0, or undefined, if they are made undefined or irrelevant by the values of other parameters hierarchically above them. Here, however, we identify the undefined with the default state. This is a simplification which may not be generally valid, and which we plan to remove in future studies. With such a simplification, each language is then assigned a binary list of syntactic parameter values. The implications already limit the number of possible combinations the latter can take. Our goal is to study how the additional random couplings further affect changes in time of the binary vector associated with a community of speakers.
In summary, we assume universal interactions $\{J_{ij\dots k}\}$ as between clusters of syntactic parameters, which are comprised in the most basic model of Implications $\{I_{ij\dots k}\}$ and of random pairwise couplings, or Weights $\{W_{ij}\}$:
$J_{ij\dots k}=W_{ij}+\zeta I_{ij\dots k},$ (1)
Here, $\zeta$ denotes the overall strength of the implications , relative to that of the dyadic interactions. Again, we use the notation $\{J_{ij\dots k}\}$ for consistency with the statistical physics literature [22].
The dyadic interaction weights $W_{ij}$ are expressed as a combination of symmetric and antisymmetric components:
$W_{ij}=\cos(\phi)W_{ij}^{S}+\sin(\phi)W_{ij}^{A},$ (2)
where both $W_{ij}^{S}=W_{ji}^{S}$ and $W_{ij}^{A}=-W_{ji}^{A}$ are taken to be random numbers drawn from the same normal distribution of zero mean and unit variance. The variable $\phi$, ranging as noted above from $\phi=0^{\circ}$ (fully symmetric weights) to $\phi=90^{\circ}$ (fully antisymmetric) is expected to play a crucial role, as symmetry generally tends to promote stability.
In analogy with statistical physics models of glassy Glauber dynamics [23] we update the values of the syntactic parameters in discrete time by choosing at each time step a random sequence (a permutation $i_{p}(i)$ of the $N$ parameters) and using in turn the zero temperature equation
$S^{\mu}_{i_{p}}(t+1)=\text{sign}\left[\sum_{j\dots k}J_{ij\dots k}S^{\mu}_{j}(t^{\prime})\dots S^{\mu}_{k}(t^{\prime})-\theta_{i}\right],$ (3)
where $S^{\mu}_{j}(t^{\prime})$ represents the state of the $j$-th syntactic parameter in language $\mu$ at time $t$, if $j$ has not yet been updated, $i_{p}(j)>i_{p}(i)$, and at time $t+1$ if already updated; and $\theta_{i}$ is a threshold controlling, in general, the likelihood that parameter $i$ would remain in the unmarked state ($S_{i}=-1$). In our current model, the threshold $\theta_{i}$ and the implication weights $I_{ij\dots k}$ are so set that, in the absence of the random couplings $W_{ij}$, the sum $\sum_{j\dots k}I_{ij\dots k}S^{\mu}_{j}(t^{\prime})\dots S^{\mu}_{k}(t^{\prime})-\theta_{i}$ is exactly zero when the other parameter values are such that parameter $i$ can choose state (hence, with equal probability), and is negative ranging downward to -1, in all other conditions. Since implications are effectively Boolean expressions, this stipulation uniquely determines both $\theta_{i}$ and $I_{ij\dots k}$.
II.1 Symmetric random couplings stabilize syntax
We begin by simulating the future evolution of each of the 58 natural languages in the database, 10 times, with each of 100 distinct random coupling sets, for a total of 1000 runs per language. We collect the number of parameters flipping at each time step and for a first statistical assessment we do not look at each language individually, but rather at averages over a total of 58,000 data points. The results are plotted in Fig. 1.
The black curve in Fig. 1 shows the number of parameters flipping over time, for $\zeta\to\infty$, i.e. subject solely to implicational constraints, those typically emerging in linguistic analyses. Syntactic structures then continue to evolve rapidly, with 20–30 parameters flipping at each time step, which is obvious considering that some of the parameters in the PCM database are not constrained by implications (13 out of 94) and therefore flip with probability 1/2, enabling other parameters to also flip in their wake (see Fig. 10 in the Appendix). In this model, therefore, were it not for the random couplings, natural languages would be fluid, and change over a time scale set, essentially, by the time step.
Introducing random symmetric connections ($\phi=0^{\circ}$) constrains these dynamics: a relatively large number of parameters (e.g., 10–30) still flip during the first or first few time steps, after which the system rapidly gets immobilized, or frozen in a configuration from which it cannot escape (Fig. 1a), regardless of the exact strength of the implications, $\zeta$.
Obviously, freezing reflects the assumption, in this particular framework, that syntactic parameters are influenced solely by endogenous compatibility, in the absence of any external inputs. Consequently, the system rapidly attains a stable equilibrium state; even very weak unknown interactions (corresponding to large $\zeta$) suffice to freeze its evolution. This scenario, albeit a limit case, is interesting insofar as it is not an automatic consequence of the assumptions in the model.
Introducing asymmetry in the random interactions, in fact, can restore fluidity to the syntax network. This is not yet evident with $\phi=30^{\circ}$ (Fig.1b) but it is for higher asymmetry values, like $\phi=60^{\circ}$, Fig.1c, all the way to full anti-symmetry, $\phi=90^{\circ}$, Fig.1d.
Therefore, the way parameters interact with the random couplings has a dramatic effect on the system dynamics: however weak they may be, what is most important is their degree of symmetry.
Partitioning syntactic parameters into classes based on their dependency order — the independent ones in class $D_{0}$, those dependent on a single other parameter in $D_{1}$, those on two, $D_{2}$, on three, $D_{3}$, and so forth — allows a better understanding of how the degree of symmetry in the random couplings affects network dynamics. The implications create a hierarchy among parameters and, when they evolve subject to implications alone, those on top of the hierarchy are completely free, while those in stepwise lower classes are progressively more constrained, and only a smaller fraction flips at every time step. Random symmetric couplings, however weak, make all parameters freeze, irrespective of their position in the hierarchy: syntax evolution is blocked, obstructed by the multitude of mutual constraints (at zero temperature, i.e., in the absence of fast noise). When the random couplings are anti-symmetric, SM Section I and Fig. S1 show that the original hierarchy is expressed again, in a quantitatively similar fashion, at least close to the deterministic limit ($\zeta$ large). Again, the fluidity of the system relies heavily on that of the 13 central $D_{0}$ parameters, and lower classes are progressively more constrained, even though the random couplings are not informed by the class structure.
Figure 2: Asymmetry makes the system fluid. The fraction of parameter vectors still changing at $t=100$ as a function of the asymmetry angle $\phi$. Data for 6 different $\zeta$ values. The dashed line indicates the presumed critical value of $\phi$, $\phi_{c}=30^{\circ}$, which is see to mark the crossing point of the curves for different values of $\zeta$. A weaker implication strength results in a more abrupt transition. Note that datapoints for $\zeta=1,0.1$ and $0.01$. as well as for $\zeta=100$ and $1000$, are superimposed onto each other, indicating that the critical $\zeta$ range is around $\zeta\simeq 10-50$.
Figure 3: Phase diagram for the basic model. The heatmap shows the mean number of parameters flipping at the $100^{th}$ time step, on a logarithmic color scale. The angle $\phi$ (in degrees) is on the x-axis, and the strength of the implications $\zeta$ is (logarithmically) on the y-axis. Towards the bottom of the diagram (small $\zeta$), implications become irrelevant, and the model behaves like a simple unstructured disordered system. For nearly symmetric interactions (up to $\phi\approx 30^{\circ}$, marked by the white boundary), dynamics come to halt, as in a spin glass, while more asymmetric interactions lead to fluid behavior. At the top of the diagram, large $\zeta$ and weak but sufficiently symmetric random interactions still stabilize parameters in metastable states of indefinite duration. By increasing asymmetry, the system transitions to a fluid state, though dynamics are slower compared to those occurring with weaker implications and similar degrees of asymmetry.
III A phase diagram for syntax change
The fluidity of the system - whether the network of syntactic parameters evolves or not - appears to be relatively independent of the natural language considered, i.e., the starting binary vector of each simulation, but also, largely, of the strength $\zeta$ of the implications. The critical determinant of fluidity is the degree of symmetry $\phi$ of the random interactions. Its role can be understood by approaching the limit $\zeta\to 0$, in which implications become irrelevant. In such a limit, our model becomes an asymmetric couplings variant of the Sherrington-Kirkpatrick model of a spin glass [22].
There have been relatively few studies of models of spin glass systems subject to asymmetric interactions [24, 25, 26], and they differ in details like the statistics of the random couplings, the update rule or the parametrization of the degree of asymmetry. The couplings in these systems, $J_{ij}$, are typically expressed as a combination of symmetric and antisymmetric components using a slightly different notation, $J_{ij}=J_{ij}^{S}+kJ_{ij}^{A},$ ranging from strictly symmetric ($k=0$) to purely antisymmetric $k\to\infty$. The symmetry factor $k$ is in turn related to the commonly used quantity $\eta$, defined as $\eta:=\langle J_{ij}J_{ji}\rangle/\langle J_{ij}^{2}\rangle,$ which leads to the relations, useful to bridge across studies, $\tan\phi=k=\sqrt{(1-\eta)/(1+\eta)}.$ Note that in these 3 alternative formulations, an average lack of either symmetry or anti-symmetry, that is, simple asymmetry, corresponds to $\phi=45^{\circ}$ or $k=1$ or $\eta=0$.
It has been shown with a variety of approaches and in a number of slightly different variants that, as the degree of coupling asymmetry increases, the system crosses over from a glassy to a fluid regime. The transition may be observed by looking at cycle length [24, 25], correlation entropy [26], or relaxation time [24, 25, 26]. With nearly symmetric couplings, for instance, the cycle length remains independent of system size, while the relaxation time scales as a power of $N$. For highly asymmetric couplings, on the other hand, both cycle length and relaxation time grow exponentially with $N$. With parallel rather than sequential updating, Nutzel and Krey have identified a subtle transition point at $\eta=0.5$, where the attractor length changes significantly [24, 25]. For $\eta>0.5$, the system gets trapped in fixed points or cycles of length 2. For $\eta<0.5$ (more asymmetric couplings), the system enters chaotic dynamics, characterized by long cycles. Crisanti et al, on the other hand, have observed that the system becomes increasingly unpredictable once the asymmetry angle exceeds a critical threshold, and have proposed a scaling law for the mean Shannon entropy around $k_{c}\simeq 0.5$ (i.e., $\eta_{c}\simeq 0.6$, or in our terms $\phi_{c}\simeq 26^{\circ}-27^{\circ}$) [26]. They note that it could be a bona fide phase transition or a simple cross-over. Which of these two slightly divergent findings is in line with our own?
To better determine numerically the point at which the transition occurs, we count how often, on average across languages, more than 3 parameter still flip from the $99^{th}$ to the $100^{th}$ time step, at which point the average number of flips per time step is roughly constant. This analysis is based on 250 simulations per language (with 50 distinct weak coupling sets, each consisting of 5 runs). Then, we extract the fraction of moving parameter vectors (i.e., those that are flipping 3 or more parameters). Fig. 2 shows that for decreasing values of $\zeta$, the transition becomes progressively sharper, around the critical value $\phi_{c}=30^{\circ}$. Weaker implications bring the network of parameters closer to its statistical physics analog, and lead to a more abrupt transition, even keeping the same low $N=94$. The syntax network therefore shows, particularly clearly for small $\zeta$, a rather sharp transition, apparently for $\phi$ close to $30^{\circ}$, in agreement with the critical point proposed by Nutzel and Krey at $\eta_{c}=0.5$ (or, equivalently, $k_{c}=1/\sqrt{3}$).
Figure 3 shows a schematic phase diagram in the $\zeta$$\phi$ plane, obtained by counting the average number of flipping parameters at the $100^{th}$ time step, i.e., in the asymptotic regime. The sharp change from a metastable to a fluid regime in the bottom part of the diagram is broadly consistent with the putative phase transition seen in asymmetric spin glass models the correspond to the limit $\zeta\to 0$ of our syntax network. As $\zeta$ increases, the crossover becomes smoother, but still remarkably well defined (considering also the small size of our system, $N=94$) and centered around the same value $\phi=30^{\circ}$. Note that along the vertical axis, the implication strength spans several orders of magnitude, on a logarithmic scale, and that a major decrease in fluidity is apparent, in the antisymmetric limit, only when $\zeta$ gets larger than $\approx 20-50$. The factor $\zeta$ modulates the strength of individual implications, but since there are fewer implicational coefficients (a total of 753, 659 multi-unit coefficients $I_{ij\dots k}$ and 94 thresholds $\theta_{i}$) than pairwise couplings ($94\times 93=8742$, in our fully connected network) and they are also weaker, only above the range $\zeta\approx 20-50$ they can significantly constrain the fluid dynamics. Irrespective of $\zeta$, for sufficiently symmetric interactions, up to $\phi=30^{\circ}$, the system exhibits glassy behavior with short-lived dynamics, and syntactic parameters rapidly settle into one of a disordered multitude of metastable states - a glassy phase. This parallels the slow dynamics and seemingly directionless diversification of grammars. The simulations then indicate that even weak yet sufficiently symmetric random interactions can stabilize multiple states. The behavior for large $\zeta$ is markedly different from the apparent asymptotic limit $\zeta\to\infty$, i.e., no random couplings, where syntactic parameters remain fluid, led by the 13 free ones among them. These findings suggest that incorporating disorder as an additional interaction term — alongside the intrinsic implications — plays a critical role in generating slow dynamics and long-term persistence in language evolution.
Figure 4: A Hebbian term is effective only if strong. Heatmap showing again the average number of parameters flipping at the $100^{th}$ time step, with fully antisymmetric random interactions ($\phi=90^{\circ}$) when adding a stabilizing Hebbian term, as described in the text. The x-axis now denotes the factor $\rho$, which regulates the strength of the Hopfield term, and the y-axis again $\zeta$. Weak Hebbian interactions (up to $\rho\approx 1$), do not affect the fluid behavior prevailing with antisymmetric random interactions, as seen in Fig. 3. As the Hebbian interaction strength increases, the network becomes frozen. The cross-over appears rapid, but note the logarithmic $\rho$ scale.
Figure 5: An apparent OR rule for stabilizing syntax. The heatmap with the number of flipping parameters at the $100^{th}$ time step, for small $\zeta=0.01$, in the $\phi-\rho$ plane. Syntactic parameters stabilize (the blue region) if at least one of two conditions hold: sufficiently symmetric random interactions (up to $\phi\approx 30^{\circ}$) or a strong Hebbian term $\rho\gtrsim 1$. As discussed below, the two conditions are not however equivalent.
IV Introducing a Hebbian term
The model considered so far ignores the evidence provided by existing natural languages, in setting the random couplings - they are assigned purely at random. It may be hypothesized, however, that if specific realizations of syntax have arisen in human history, this may reflect interactions about particular syntactic parameters as they are implemented in the human brain. One simple-minded approach to delve into this complex and ill-defined issue is to introduce a set of Hebbian autoassociative synaptic weights, which in the syntax model would simply be additional pairwise interaction terms, strictly symmetric, that tend to stabilize existing realizations of language parameters. Inspired by the Hopfield model of an autoassociative network [27], we may identify the 58 languages in the PCM database as the ‘memories’ of the syntactic network. The new weights, denoted as $W_{ij}^{\text{Hop}}$, to be included in the interaction term in Eq. 1, $J_{ij\dots k}=W_{ij}^{\text{RND}}+\zeta I_{ij\dots k}+\rho W_{ij}^{\text{Hop}},$ are written as
$W_{ij}^{\text{Hop}}=1/N\sum_{\mu}S^{\mu}_{i}S^{\mu}_{j},$
where the factor $\rho$ regulates their overall strength, $S^{\mu}_{i}$ represents the binary state of syntactic parameter $i$ for the $\mu^{th}$ language, and $\mu=1,\ldots,L$ samples the available database.
Figure 4 presents a schematic phase diagram in the $\zeta$-$\rho$ plane for a maximally fluid system – one with fully antisymmetric random interactions ($\phi=90^{\circ}$) – in the asymptotic regime (at the $100^{th}$ time step). Whatever the strength of the implications, there is clearly a cross-over from a fluid to a stable regime once the Hebbian term becomes effective, $\rho\gtrsim 1$: the Hebbian term then stabilizes the system even when the other interactions are strictly antisymmetric.
Figure 5 shows a similar phase diagram in the $\phi$-$\rho$ plane, when implications are negligible, $\zeta=0.01$ (note that $\rho$ has been moved to the y-axis). It appears to demonstrate that the stable regime, when syntactic parameters have completely stopped evolving by the $100^{th}$ time step, can be realized with either of two conditions, or both: when the Hebbian term is strong, $\rho>1$, or random interactions are sufficiently symmetric, $\phi<30^{\circ}$. The regime boundaries appear surprisingly close to a Boolean rule.
In the high-$\rho$ regime, however, all syntactic configurations rapidly converge to a unique stable state, as shown in Fig.6, which maps the average Hamming distance between vectors evolving from different starting configurations.
The reason for convergence to a single language, rather than to one of a disordered multitude of metastable states, is that unlike disorder, which halts parameter evolution through frustration, the Hopfield term, a form of order, suppresses linguistic diversity. The Hebbian term, therefore, does not seem able to account for the observed multiplicity of languages, when it dominates, and it does not lead to observable effects when it does not dominate. It could be, it should be noted, that a weighted contribution by the languages in the dataset, or the use of a wider sample of different languages, might each prevent the collapse and preserve multiplicity, i.e., language diversity. The available evidence is however insufficient to structure such differential weights.
Figure 6: Different languages rapidly converge, for $\rho>1$. Mean pairwise Hamming distances between realizations of distinct languages subject to identical random couplings, after only two time steps (($t=2$, panel (a)), and in the asymptotic regime ($t=100$, panel (b)). These distances vanish when the Hebbian term is dominant ($\rho>1$), indicating rapid collapse, already after 2 time steps, onto a single syntactic structure. Note that in the fluid regime, bottom right corners, Hamming distances $\sim 47$ imply that typically half the parameters agree, between any two vectors, and half do not.
In the remainder of this study, therefore, the Hebbian term will not be considered further, aside from a detail in Fig. 9 and for the analysis of syntactic distances presented in the Supplementary Material.
Figure 7: Languages stabilize rapidly. Hamming distance between realizations at the final time step ($t=100$) and those at earlier stages ($t=2$, $10$, and $40$) for Italian syntax, as an example. While such distance remains substantial in the region of high asymmetry to the right (larger when implications are weaker, bottom right), in prevailingly symmetric region in the left the syntactic vector has changed, at $t=100$, only if compared to very early stages of evolution, such as $t=2$, (a). For $\phi$ close to 0 it has already ceased to evolve by $t=10$, (b), and definitely by $t=40$, (c), all the way to nearly $\phi\sim 30^{\circ}$.
Figure 8: Evolution of inter-language and intra-language distances. The diagrams, in the $\zeta$$\phi$ plane (at $\rho=0$), show both at early ($t=2$, top row) and asymptotic times ($t=100$, bottom row) the average Hamming distance between simulations starting from the same natural language (distance in condition S, left) and different natural languages (distance in condition D, middle), as well as their difference, (D-S, right), when using always identical random couplings. With time, both average distances diminish in areas characterized by strong implicational dynamics, especially when asymmetry is minimal. Still, they do not vanish: the average Hamming distance D between descendants of different languages tends (asymptotically, or for $t=100$) to decrease from D $\sim 35-37$ for weak implications ($\zeta<1$) thorough an intermediate range (e.g., D $\sim 18$ for $\zeta=10$, the star symbol) down to values D $\sim 3.1$ when implications dominate, for $\zeta>100$ (round symbol). Their average difference D-S illustrates the gradual fading of the phylogenetic/historical signal, i.e., the trace of shared ancestry among languages, everywhere except in the nearly symmetric and low (or at most intermediate) implicational strength regime, bottom left. Even there, the signal is however weak, a difference in at most $4-5$ parameters, compared to the dispersion observed starting from the same initial condition (left column).
V Preservation of a phylogenetic signal
Tracing the persistence of syntactic structures in language evolution requires quantifying in some way how they diverge over time. To this end, we use as a first approximation the same simple notion of distance between languages used above to assess stability, i.e., the Hamming distance between different syntactic vectors - the number of parameters where the two vectors differ, ranging from 0 to 94 - where we do not rank syntactic parameters in terms of importance, or along their hierarchy. This metric ignores a wealth of linguistic knowledge, but can give a first impression, within the limits of our model, as to the long-term divergence or convergence among instantiations of syntax. We note that very similar observations can be based on the use of the Euclidean distance between vectors.
As an example of temporal evolution of a syntactic system within our model, we compare the presumed asymptotic state at $t=100$ with selected intermediate stages ($t=2,10,40$) for the case of a real-world starting point — Italian — in Fig. 7. When implications are irrelevant and asymmetry is high, the vector at $t=100$ remains very distant from its ancestors at essentially all preceding $t$ stages – note that a distance $\sim 47$ implies that half the 94 parameters take the same value in the two configurations, half a different value: effectively no syntax correlation. In other words, parameter values wander randomly in syntactic space, and reach no stable state nor do they approach an asymptotic state. Something similar occurs, albeit with reduced fluidity, when implications are stronger, but random couplings are still asymmetric. In contrast, the distance from the asymptotic configuration gradually vanishes if random couplings are prevailingly symmetric, and in fact when they are close to full symmetry syntactic parameters are already stable after a few time steps. These results are consistent across different languages.
If that is the time course of simulations starting from the same point, we need to compare them with those starting from other initial syntax vectors. A phylogenetic signal encodes the relatedness among descendants of the same syntactic configuration that persists, despite the initially rambling trajectories in syntactic space, produced by the random couplings. We obviously need to keep the assignment of random couplings the same, and compare distances between different runs with the Same starting point (case S), and with Different starting points (case D). Figure 8 presents this comparison, as a grand average, in the entire $\zeta$-$\phi$ plane (at $\rho=0$), for configurations evaluated at two distinct time points: $t=2$ (top row) and $t=100$ (bottom row), corresponding to the early and asymptotic phases of the model, respectively. In each case, the main factor affecting these distances, which we label simply S and D like the cases they refer to, is the strength of the implications: when they are strong, they constrain the dynamics and both types of distance grow less - in fact, they decrease, reaching lower values at $t=100$ than at $t=2$. Thus, while for minimal asymmetry and in the low implication region, $\zeta<1$, the average D $\gtrsim 35$, in the strong implication regime, $\zeta>100$, it goes down below one tenth of this value, D $\sim 3.1$. When implication compete with the random couplings, $1001<\zeta<$ 100, one observes intermediate values; but, whatever $\zeta$, there is always dispersal, i.e., a multiplicity of trajectories, which do not converge to a single language.
Interestingly, however, when we take the difference S-D, the only region where it remains above zero at $t=100$ (even though with small values) is the region of nearly symmetric random couplings and weak or intermediate implications, up to about $\zeta\approx 50$. Strong implications, $\zeta>100$, even with symmetric random couplings, retain a phylogenetic signal only for the first few time steps (top left region in the top right panel of Fig. 8). This indicates that in this model the phylogenetic signal is washed away, after a sufficient number of time steps, if implications are strong.
Figure 9: Temporal dynamics of the phylogenetic signal. (a) The time course of the phylogenetic signal with $\phi$ fixed at $15^{\circ}$ and $\rho=0$, showing the rapid decrease in signal followed by stabilization at a steady-state value. A robust asymptotic signal is still observed for $\zeta=1$, i.e., when implications and random contributions are balanced. As $\zeta$ increases, indicating a greater influence of implications over random terms, the steady-state signal decreases. (b) The impact of varying $\rho$ and $\phi$, while keeping $\zeta=10$ constant. With limited asymmetry, $\phi=15^{\circ}$, a small Hebbian term (the red dashed line for $\rho=0.1$) leads the system to approximate the steady-state signal faster (compare to the yellow line for $\rho=0$). Further asymmetry, i.e., increasing $\phi$ leads to the vanishing of the signal, concurrent with the cross over to the fluid regime.
To further detail the time course of the average phylogenetic signal in the model, Figure 9 shows it as a function of $\zeta$ for $\phi=15^{\circ}$, in a, and as a function of $\phi$ for $\zeta=10$, in b. The signal rapidly decreases over a few iterations and then stabilizes at a steady-state value. This steady-state level depends on the variables $\phi$ and $\zeta$, and is stronger when they are both low. Notice in b that a weak additional Hebbian term, insufficient to lead to the collapse into a single stable state, does not alter the phylogenetic signal either, aside from initially accelerating the approach to the asymptotic state. Notice also that already at $\phi=30^{\circ}$ the signal has disappeared, as the syntactic network has crossed over into the fluid regime.
VI Conclusions
This study presents a framework for modeling syntactic evolution through the lens of disordered systems, bridging formal linguistics and statistical physics. By treating syntactic parameters as binary variables subject to both logical and typological implications and to random interactions, the model aims to capture the dual behavior of syntactic systems: their remarkable persistence across time and their capacity for gradual yet unpredictable diversification.
Our results indeed indicate that in a portion of its phase space the network exhibits glassy dynamics, slow and gradually diverging, similar to that of spin glasses. Whether in the $N\to\infty$ limit the glassy regime would become a bona fide spin glass phase remains to be seen, but the glassy behavior is evident already in the network of limited size that models the $N=94$ syntactic parameters of the PCM database. Thus, the model suggests that syntactic evolution may not merely be the product of deterministic or functionally driven processes, and might instead be shaped by subtle, usually unobservable interactions that likely arise from the way syntax is realized in the brain of the speakers. The critical asymmetry value $\phi_{c}=30^{\circ}$, presumably marking a phase transition in the similar statistical physics model, can serve as a valid reference point for the syntax model as well, as it is not much affected by other factors, in particular the strength $\zeta$ of the implications. The suggestion from the model, then, is that natural language parameters interact, perhaps due to their implementation in the brain, via couplings that are, if not strictly symmetric, at least prevailingly symmetric (quantitatively, $\phi<30^{\circ}$).
Of course, in our idealized framework – no external inputs whatsoever and zero ‘temperature’, i.e., zero noise – syntactic systems can become ‘frozen’ in certain configurations, while in more realistic scenarios they would continue to evolve, though slowly, in the corresponding region of the phase diagram. It should be noted that a separate study has shown that a disordered network of binary variables interacting randomly with each other as well as with a set of multi-state variables (also randomly) can be even glassier, that is slower, than if it were isolated [28]. The suggestion would then be that if syntax – the binary parameters – interacts randomly with e.g. the lexicon, or phonology, or semantics – which can all be formulated in terms of multi-state variables – it might evolve even slower than on its own.
The introduction of a Hebbian memory term would stabilize syntactic configurations, even with asymmetric random couplings. However, an effective Hebbian term ($\rho>1$) would lead all languages to collapse onto the same syntax, at least in our implementation of the Hebbian term. More sophisticated versions of the same idea might avoid the collapse. For example, it might be argued that treating closely related languages, like several of those from the Indo-European family that comprises nearly half of the current PCM database, as adding their separate Hebbian contributions might be excessive, theoretically dubious, and might be the key factor leading to the collapse. Still, it is not obvious at this stage how to construct a ‘balanced’ Hebbian term, mimicking the effect of orthogonal memories in autoassociative memory networks [27].
Understanding better the conditions favoring glassy dynamics is also made urgent by the observation that in the present model, even in the absence of a Hebbian term, the phylogenetic signal rapidly vanishes if implications are sufficiently strong to really act as constraints, $\zeta\gtrsim 10-50$ (see Fig. 8); therefore invalidating the original assumption that adding weak interactions $W_{ij}$ to strong implications would suffice to lead to both divergence (the multiplicity of metastable states) and persistence (the phylogenetic signal). In the model as presented here, even with no Hebbian term ($\rho=0$), and with prevailingly symmetric random couplings ($\phi<30^{\circ}$, better if $\phi\sim 0^{\circ}$) divergence is maintained, to variable degrees, for all values of $\zeta$, but persistence requires weak or at most intermediate implications $\zeta\lesssim 50$.
This study, in conclusion, aims to present a first version of the mapping of syntactic parameters onto a recurrent network model, and needs to be extended in several directions before it can lead to well-grounded inferences. One direction we see as critical is a refined modeling of the undefined state of parameters, the definition of which has been disabled by the values taken by other parameters. Further, we should consider distinguishing different parameters, depending on the nature of their putative default state [29]. We plan to pursue these directions in future work.
Acknowledgments
This study was started with an important initial contribution by Kwang Il Ryom, then a PhD student at SISSA, who should have coauthored it, had it been possible to contact him in recent months. We are also indebted to the many researchers who have contributed to the PCM database, foremost Cristina Guardiano and Paola Crisma; and to Eric De Giuli for stimulating discussions.
Data Availability
The PCM Database has been published, e.g. in [9] and expanded versions are likely to be published in the near future. The software used in the present study will be made freely available upon publication.
Author Contributions
H.Y. carried out the numerical experiments and drafted the manuscript; G.L. conceptualized the study, structured the syntactic implications in the PCM database and revised the manuscript; A.T. conceptualized the study, coded the implications as polysynaptic weights and drafted the manuscript.
Appendix A Dynamical roles of syntactic dependence classes
To clarify how implications constrain changes in the values of the syntactic parameters, it is useful to group them into classes according to their dependence depth — that is, the number of other parameters their relevance depends upon — labeling these groups $D_{0},D_{1},D_{2},\ldots,D_{6}$. For example, a parameter in $D_{3}$ is only relevant if 3 other parameters take specific values; otherwise it is undefined, which in our model is conflated with taking the default value -1.
Figure 10: Dynamical evolution of syntactic parameters grouped by dependence depth. (a) Temporal evolution of the normalized number of flipped parameters for each dependence class $D_{0},D_{1},\ldots,D_{6}$ in the deterministic limit ($\zeta\to\infty$). The 13 independent parameters in class $D_{0}$, which are unconstrained by implications, are the most active, flipping around half the times. (b) Same as (a), but with added weak symmetric random interactions ($\zeta=1000$, $\phi=0^{\circ}$). They suffice to freeze the system of parameters. (c) Same as (a), but with weak antisymmetric stochastic interactions ($\zeta=1000$, $\phi=90^{\circ}$), which keep the system fluid.
First, we look at network dynamics in the deterministic limit ($\zeta\to\infty$), where dynamics are governed purely by implications, in the absence of the extra interactions, which in our model are randomly assigned. Figure 10a shows the temporal evolution of the normalized count of flipping parameters for each syntactic class. Independent parameters, the core set of 13 syntactic features in $D_{0}$ not dependent on implications, of course flip every second time step, on average. Changes in these free parameters propagate through the hierarchy, making parameters in $D_{1}$ those with the next highest probability to flip, then those in $D_{2}$, and so on. Ultimately, therefore, all syntactic changes are enabled, in this limit, by flipping the 13 independent parameters, out of the total 94, in the absence of random couplings. When the latter are introduced, even with minimal strength ($\zeta=1000$, see Fig. 10a and b for $\phi=0^{\circ}$ and $\phi=90^{\circ}$, respectively), the free parameters are not free anymore: for them, unconstrained by implications, the random couplings are the interactions there are, and they suffice to freeze the evolution of the free parameters in a few time steps, if symmetric or nearly symmetric. The other parameters then freeze as a result, see Fig. 10a for $\phi=0^{\circ}$. Beyond approximately $\phi=30^{\circ}$ the random interactions are asymmetric enough to keep the system fluid, and the fluidity is expressed in a similar hierarchy of flipping rates as the one seen in the absence of random couplings (as Fig. 10c shows for $\phi=90^{\circ}$).